How many gallons of water is needed to add to a 1 gallon 100% concentration of anti freeze to form a mixture of 30%?
Answer by ewatrrr(24785)

Hi,
30% states how much of the TOTAL mixture will be anti freeze!
30% = (30/100) = .30
.
Let T be the amount of the TOTAL mixture: Problem States there will be 1 gallon of antifreeze in the mixture
.
.30*T = 1gallon
.
T = 3 1/3 Gallons ( 1 Gallon of which is antifreeze)
.
Need to add 2 1/3 gallons of water to make the proper mixture